# Denominators in Lusztig’s asymptotic Hecke algebra via the Plancherel formula

### v5

@@ 3.2 The functions $f_{w}$ for general $\mathbf{G}$
Let $w\in\tilde{W}$. Then $f_{w}(1)$ is a rational function of $q$ with poles drawn from a finite set of roots of unity depending only on $\tilde{W}$. The numerator is a Laurent polynomial in $q^{1/2}$. The denominator depends only on the two-sided cell containing $w$. If $w$ is in the lowest two-sided cell, then the denominator divides the Poincaré polynomial of $\mathbf{G}$.
@@
More generally, let $w\in\tilde{W}$ and $M_{P}$ be a Levi subgroup corresponding to the two-sided cell containing $w$. Let $N=\operatorname{rk}A_{P}$. Given a coroot $\alpha^{\vee}$ and a basis of the cocharacter lattice as explained in Section 2.7, we write $\alpha^{\vee}=z_{1}^{e_{1}}\cdots z_{n}^{e_{n}}$ for integers $e_{i}=e_{i}(\alpha)$.
@@ 3.3 Relating $t_{w}$ and $f_{w}$
[…]
@@ 3.3.1 Completions of $\mathbf{H}$ and $J\otimes_{\mathbb{C}}\mathcal{A}$
[…]
@@
[…]
@@
[…]
@@ 3.3.2 The functions $f_{w}$ and the basis elements $t_{w}$
[…]
@@
[…]
@@
Let $w\in\tilde{W}$. Then the degree in $q$ of the numerator of $f_{w}(1)$ is bounded uniformly in $w$ by some $N$ depending only on $\tilde{W}$, and hence in the notation of Proposition 2 above, we have $a_{1,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $w\in\tilde{W}$.
@@
Now, the denominator of $f_{w}(1)$, and hence $\deg d_{1}$, depends only on the two-sided cell containing $w$, again by Corollary 2 and Theorem 12. We can also bound $\deg h_{1}$ uniformly in terms of $\tilde{W}$, as it depends only on the Plancherel measure and the finitely many possible formal degrees appearing in the parametrization of the $I$-spherical part of the tempered dual of $G$. This proves the lemma. ∎
@@
[…]
By Corollary 2 and Theorem 12 , there is a formal power series in $\hat{\mathcal{A}}^{-}$ with constant term equal to $1$ that specializes to the denominator of $A_{1,w}$ when $\mathbf{q}=q$. Moreover, there is a unique formal Laurent series $a_{1,w}\in\hat{\mathcal{A}}$ such that $a_{1,w}(q)=A_{1,w}$ for all prime powers. Indeed, $a_{1,w}$ is convergent for $\mathbf{q}=q$, and the difference of any two such series defines a meromorphic function of $\mathbf{q}^{-1/2}$ outside the unit disk with zeros at $q=p^{r}$ for every $r\in\mathbb{N}$. As these prime powers accumulate at $\infty$, such a meromorphic function must be identically zero.
@@
By definition, $f_{w}(x)=A_{x,w}$. On the other hand, according to Lemma 2, we have
$q^{-\ell(x)}(f_{w}\star T_{x^{-1}})(1)$ $=q^{-\ell(x)}\left(\tilde{\phi}(t_{w})\star\tilde{\phi}\left(\phi_{q}(T_{x^{-1}}^{\dagger})\right)\right)(1)$
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}(T_{x^{-1}}^{\dagger})\right)(1)$
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})-\ell(y)}Q_{y,x^{-1}}(q)(C^{\prime}_{y})^{\dagger}\right)\right)(1)$ (24)
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)(C_{y})\right)\right)(1)$ (25)
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\\ d\in\mathcal{D}\\ a(d)=a(r)\end{subarray}}h_{y,d,r}t_{r}\right)(1)$ (26)
$=q^{-\ell(x)}\tilde{\phi}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ d\in\mathcal{D}\\ a(d)=a(r)=a(w)\end{subarray}}h_{y,d,r}t_{w}t_{r}\right)(1)$
$=q^{-\ell(x)}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ a(r)=a(w)\end{subarray}}h_{y,d_{w},r}(f_{w}\star f_{r})(1),$ (27)
@@
In line (24), we rewrote $T_{x^{-1}}$ in terms of the $C^{\prime}$-basis of $H$, using the inverse Kazhdan-Lusztig polynomials $Q_{y,x^{-1}}$. In line (25), we applied the Goldman involution (see Lemma 1). In line (26) we applied Lusztig’s map $\phi_{q}$, and then in line (27), we applied the map $\tilde{\phi}$. Also in line (27), we used that left (respectively right) cells give left (respectively right) ideals of $J$, and so $t_{w}t_{r}$ is an integral linear combination of $t_{z^{-1}}$, with $d\sim_{L}z^{-1}\sim_{R}d_{w}$. By lemma 9, $f_{z^{-1}}(1)\neq 0$ only if $d=d_{w}$, that is, $z^{-1}\sim_{L}d_{w}$.
@@
as $\ell(x)\to\infty$. This completes the definition of $\phi_{1}$ as a map of $\mathcal{A}$-modules.
It is easy to see that $\phi_{1}$ is a morphism of rings, essentially because $\tilde{\phi}$ is. Indeed, we have
$\phi_{1}(t_{w}t_{w^{\prime}})=\sum_{z}\gamma_{w,w^{\prime},z^{-1}}\sum_{x}a_{x,z}T_{x}$ (29)
@@
$\phi_{1}(t_{w})\phi_{1}(t_{w^{\prime}})=\sum_{x}a_{x,w}T_{x}\cdot\sum_{y}a_{y,w^{\prime}}T_{y}$ (30)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Note that as $\tilde{\phi}\circ\phi_{q}=()^{\dagger}$ on $H$ for all $q$, we have that $\phi_{1}=()^{\dagger}\circ\phi^{-1}$ on $\phi(\mathbf{H})$. This says that the diagram commutes.
[…]
$\phi_{1}\left(\sum_{w}b_{w}t_{w}\right)=\sum_{y}b^{\prime}_{y}T_{y},$
@@
We must first show that this infinite sum of elements of $\hat{\mathcal{A}}$ is well-defined. By Lemma 8 and (27), we have that there is $M\in\mathbb{N}$ such that $a_{y,w}\in(\mathbf{q}^{1/2})^{M}\hat{\mathcal{A}}^{-}$ for all $y,w$. Therefore $b^{\prime}_{y}$ is well-defined, and as $a_{y,w}\to 0$ as $\ell(y)\to\infty$, we have $b^{\prime}_{y}\to 0$ as $\ell(y)\to\infty$. Therefore $\phi_{1}$ extends to $\mathbfcal{J}$.
To show continuity, it suffices to show that if $\{\sum_{w}b_{w,n}t_{w}\}_{n}$ is a sequence of elements of $\mathbfcal{J}$ tending to 00 as $n\to\infty$, then
$\sum_{w}b_{w,n}\phi_{1}(t_{w})=\sum_{y}b^{\prime}_{y,n}T_{y}\to 0$
[…]
@@
$(\phi\circ()^{\dagger})^{-1}(t_{1})=f_{1}(1)\sum_{w\in\tilde{W}}(-1)^{\ell(w)}q^{-\ell(w)}T_{w}.$
@@
If $w=1$, everything in (27) reduces to $r=d_{w}=y=1$, and we need only recall that $Q_{1,x}(\mathbf{q})=1$ for any $x\in\tilde{W}$. This follows from unicity of the inverse Kazhdan-Lusztig polynomials and the identity $\sum_{x\leq w}(-1)^{\ell(x)}P_{x,w}(\mathbf{q})=0$ for $w\neq 1$ [BjornerBrenti, Exercise 5.17]. ∎
@@
Let $q>1$ and let $j\neq 0$ be an element of $J$. We must show that $\tilde{\phi}(j)\neq 0$. By injectivity of $\phi_{1}$, we have $\phi_{1}(j)\neq 0$. By definition of the map $\phi_{1}$, this means that there exists $q_{0}>1$, $u,s\in G^{\vee}$ such that $us=su$ with $s$ compact, and a representation $\rho$ of $\pi_{0}(Z_{G^{\vee}}(u,s))$ such that $jK(u,s,\rho,q_{0})\neq 0$. But then $jK(u,s,\rho,q)\neq 0$, $K(u,s,\rho,q)$ being a different specialization of the restriction of the same $J$-module $E(u,s,\rho)$ as for $K(u,s,\rho,q_{0})$, and $K(u,s,\rho,q)$ is also tempered. It follows that $\tilde{\phi}(j)\neq 0$. ∎
3.4 Proof of Theorem 13 for $\mathrm{GL}_{n}$

### v6

@@ 3.2 The functions $f_{w}$ for general $\mathbf{G}$
Let $w\in W_{\mathrm{aff}}$. Then $f_{w}(1)$ is a rational function of $q$ with poles drawn from a finite set of roots of unity depending only on $W_{\mathrm{aff}}$. The numerator is a Laurent polynomial in $q^{1/2}$. The denominator depends only on the two-sided cell containing $w$. If $w$ is in the lowest two-sided cell, then the denominator divides the Poincaré polynomial of $\mathbf{G}$.
@@
More generally, let $w\in W_{\mathrm{aff}}$ and $M_{P}$ be a Levi subgroup corresponding to the two-sided cell containing $w$. Let $N=\operatorname{rk}A_{P}$. Given a coroot $\alpha^{\vee}$ and a basis of the cocharacter lattice as explained in Section 2.7, we write $\alpha^{\vee}=z_{1}^{e_{1}}\cdots z_{n}^{e_{n}}$ for integers $e_{i}=e_{i}(\alpha)$.
@@ 3.3 Relating $t_{w}$ and $f_{w}$
We will now relate the Schwartz functions $f_{w}$ on $G$ to the elements $\phi^{-1}(t_{w})$ of a completion $\mathbfcal{H}^{-}$ of $\mathbf{H}$, whose definition we will now recall. In this section $\mathbf{G}$ is general.
@@ 3.3.1 Completions of $\mathbf{H}$ and $J\otimes_{\mathbb{C}}\mathcal{A}$
Let $\hat{\mathcal{A}}=\mathbb{C}((\mathbf{q}^{-1/2}))$ and $\hat{\mathcal{A}}^{-}=\mathbb{C}[[\mathbf{q}^{-1/2}]]$. Write $\mathbfcal{H}^{-}$ for the $\hat{\mathcal{A}}-$algebra
@@
In [affineII], Lusztig shows that $\phi$ extends to an isomorphism of $\hat{\mathcal{A}}-$algebras $\mathbfcal{H}^{-}_{C}\to\mathbfcal{J}$. In this way the elements $t_{w}\in J\subset\mathbfcal{J}$ may be identified with elements of $\mathbfcal{H}^{-}_{C}$ via $\phi$.
@@
We have $\mathbfcal{H}^{-}_{C^{\prime}}\subset\mathbfcal{H}^{-}$. The inclusion is continuous.
@@ 3.3.2 The functions $f_{w}$ and the basis elements $t_{w}$
We shall now explain how the map $\tilde{\phi}\colon J\to\mathcal{C}(G)^{I}$ induces a map of $\mathcal{A}$-algebras $\hat{\phi}\colon J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$.
@@
There is a map of $\mathcal{A}$-algebras $\hat{\phi}\colon J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$ such that if $\hat{\phi}(t_{w})=\sum_{x}a_{x,w}T_{x}$, then $a_{x,w}(q)=f_{w}(x)$. Moreover, there is a constant $N$ depending only on $W_{\mathrm{aff}}$ such that $a_{x,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $x,w\in W_{\mathrm{aff}}$.
@@
Let $w\in W_{\mathrm{aff}}$. Then the degree in $q$ of the numerator of $f_{w}(1)$ is bounded uniformly in $w$ by some $N$ depending only on $W_{\mathrm{aff}}$, and hence in the notation of Proposition 2 above, we have $a_{1,w}\in(\mathbf{q}^{1/2})^{N}\hat{\mathcal{A}}^{-}$ for all $w\in W_{\mathrm{aff}}$.
@@
Now, the denominator of $f_{w}(1)$, and hence $\deg d_{1}$, depends only on the two-sided cell containing $w$, again by Corollary 3 and Theorem 12. We can also bound $\deg h_{1}$ uniformly in terms of $W_{\mathrm{aff}}$, as it depends only on the Plancherel measure and the finitely many possible formal degrees appearing in the parametrization of the $I$-spherical part of the tempered dual of $G$. This proves the lemma. ∎
@@
Let $w\in W_{\mathrm{aff}}$. Write $\tilde{\phi}(t_{w})=f_{w}=\sum_{x}A_{x,w}T_{x}$ as Schwartz functions on $G$ so that $A_{x,w}=f_{w}(x)$. We must show that there is a unique element $a_{x,w}\in\hat{\mathcal{A}}$ such that $a_{x,w}(q)=A_{x,w}$ as complex numbers. We will then check that $a_{x,w}\to 0$ rapidly enough as $\ell(x)\to\infty$ for $\sum_{x}a_{x,w}T_{x}$ to define an element of $\mathbfcal{H}^{-}$.
By Corollary 3 and Theorem 12 , there is a formal power series in $\hat{\mathcal{A}}^{-}$ with constant term equal to 11 that specializes to the denominator of $A_{1,w}$ when $\mathbf{q}=q$. Moreover, there is a unique formal Laurent series $a_{1,w}\in\hat{\mathcal{A}}$ such that $a_{1,w}(q)=A_{1,w}$ for all prime powers. Indeed, $a_{1,w}$ is convergent for $\mathbf{q}=q$, and the difference of any two such series defines a meromorphic function of $\mathbf{q}^{-1/2}$ outside the unit disk with zeros at $q=p^{r}$ for every $r\in\mathbb{N}$. As these prime powers accumulate at $\infty$, such a meromorphic function must be identically zero.
@@
By definition, $f_{w}(x)=A_{x,w}$.
On the other hand, according to Lemma 3, we have, for $\omega(x^{-1})_{f}$ as defined above Lemma 2,
$q^{-\ell(x)}(f_{w}\star T_{x^{-1}})(1)$ $=q^{-\ell(x)}\left(\tilde{\phi}(t_{w})\star\tilde{\phi}\left(\phi_{q}({}^{\dagger}T_{x^{-1}})\right)\right)(1)$
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}({}^{\dagger}T_{x^{-1}})\right)(1)$
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})-\ell(y)}Q_{y,x^{-1}}(q){}^{\dagger}C^{\prime}_{y}\right)\right)(1)$ (26)
$=q^{-\ell(x)}\tilde{\phi}\left(t_{w}\phi_{q}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}(-1)^{\ell(x^{-1})}(-1)^{\ell(\omega(x^{-1})_{f})}Q_{y,x^{-1}}(q)(C_{y})\right)\right)(1)$ (27)
$=(-1)^{\ell(x^{-1})}(-1)^{\ell(\omega(x^{-1})_{f})}q^{-\ell(x)}\tilde{\phi}\left(t_{w}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\\ d\in\mathcal{D}\\ a(d)=a(r)\end{subarray}}h_{y,d,r}t_{r}\right)(1)$ (28)
$=(-1)^{\ell(x^{-1})+\ell(\omega(x^{-1})_{f})}q^{-\ell(x)}\tilde{\phi}\left(\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d\\ d\in\mathcal{D}\\ a(d)=a(r)=a(w)\end{subarray}}h_{y,d,r}t_{w}t_{r}\right)(1)$
$=(-1)^{\ell(x^{-1})+\ell(\omega(x^{-1})_{f})}q^{-\ell(x)}\sum_{y\leq x^{-1}}q^{\frac{\ell(y)}{2}}Q_{y,x^{-1}}(q)\sum_{\begin{subarray}{c}r\sim_{L}d_{w}\\ a(r)=a(w)\end{subarray}}h_{y,d_{w},r}(f_{w}\star f_{r})(1),$ (29)
@@
In line (26), we rewrote $T_{x^{-1}}$ in terms of the $C^{\prime}$-basis of $H$, using the inverse Kazhdan-Lusztig polynomials $Q_{y,x^{-1}}$. In line (27), we applied the involution $^{\dagger}(-)$ (see Lemma 2). In line (28) we applied Lusztig’s map $\phi_{q}$, and then in line (29), we applied the map $\tilde{\phi}$. Also in line (29), we used that left (respectively right) cells give left (respectively right) ideals of $J$, and so $t_{w}t_{r}$ is an integral linear combination of $t_{z^{-1}}$, with $d\sim_{L}z^{-1}\sim_{R}d_{w}$. By lemma 10, $f_{z^{-1}}(1)\neq 0$ only if $d=d_{w}$, that is, $z^{-1}\sim_{L}d_{w}$.
@@
as $\ell(x)\to\infty$. This completes the definition of $\hat{\phi}$ as a map of $\mathcal{A}$-modules.
It is easy to see that $\hat{\phi}$ is a morphism of rings, essentially because $\tilde{\phi}$ is. Indeed, we have
$\hat{\phi}(t_{w}t_{w^{\prime}})=\sum_{z}\gamma_{w,w^{\prime},z^{-1}}\sum_{x}a_{x,z}T_{x}$ (31)
@@
$\hat{\phi}(t_{w})\hat{\phi}(t_{w^{\prime}})=\sum_{x}a_{x,w}T_{x}\cdot\sum_{y}a_{y,w^{\prime}}T_{y}$ (32)
@@
Hence for infinitely many prime powers we have that the specializations of (31) agrees with those of (32), and hence (31) is equal to (32) in $\mathbfcal{H}^{-}$. A similar argument shows that $\hat{\phi}$ preserves units. ∎
@@
The proof, specifically (30), gives a necessary condition for an element of $\mathbfcal{H}^{-}$ to belong to the image of $\hat{\phi}$: the coefficients must decay asymptotically at least as fast as $\mathbf{q}^{-\frac{\ell(x)}{2}}$.
@@
$\mathbf{H}$$J\otimes_{\mathbb{C}}\mathcal{A}$$\mathbfcal{H}^{-}$$\mathbfcal{H}^{-}_{C}$$\mathbfcal{H}^{-}_{C^{\prime}},$$\scriptstyle{\phi}$$\scriptstyle{\hat{\phi}}$$\scriptstyle{\phi^{-1}}$$\scriptstyle{{}^{\dagger}(-)}$
and we have $\hat{\phi}={}^{\dagger}(-)\circ\phi^{-1}$ as morphisms of $\mathcal{A}$-algebras $J\otimes_{\mathbb{C}}\mathcal{A}\to\mathbfcal{H}^{-}$. In particular, $a_{x,w}$ has integer coefficients for all $x,w\in W_{\mathrm{aff}}$.
@@
The second claim follows from the first if we show that $\hat{\phi}$ extends to a continuous morphism $\mathbfcal{J}\to\mathbfcal{H}^{-}$, by density of $\phi(\mathbf{H})$ in $\mathbfcal{J}\simeq\mathbfcal{H}^{-}_{C}$, and the third claim follows from the second and the fact that the completions we consider are actually defined over $\mathbb{Z}$ [affineII, Thm. 2.8].
Note that as $\tilde{\phi}\circ\phi_{q}={}^{\dagger}(-)$ on $H$ for all $q$, we have that $\hat{\phi}={}^{\dagger}(-)\circ\phi^{-1}$ on $\phi(\mathbf{H})$. This says that the diagram commutes.
We now show that $\hat{\phi}$ extends to a continuous map $\mathbfcal{J}\to\mathbfcal{H}^{-}_{C}$. Let $\sum_{w}b_{w}t_{w}$ define an element of $\mathbfcal{J}$ and define
$\hat{\phi}\left(\sum_{w}b_{w}t_{w}\right)=\sum_{y}b^{\prime}_{y}T_{y},$
@@
We must first show that this infinite sum of elements of $\hat{\mathcal{A}}$ is well-defined. By Lemma 9 and (29), we have that there is $M\in\mathbb{N}$ such that $a_{y,w}\in(\mathbf{q}^{1/2})^{M}\hat{\mathcal{A}}^{-}$ for all $y,w$. Therefore $b^{\prime}_{y}$ is well-defined, and as $a_{y,w}\to 0$ as $\ell(y)\to\infty$, we have $b^{\prime}_{y}\to 0$ as $\ell(y)\to\infty$. Therefore $\hat{\phi}$ extends to $\mathbfcal{J}$.
To show continuity, it suffices to show that if $\{\sum_{w}b_{w,n}t_{w}\}_{n}$ is a sequence of elements of $\mathbfcal{J}$ tending to 0 as $n\to\infty$, then
$\sum_{w}b_{w,n}\hat{\phi}(t_{w})=\sum_{y}b^{\prime}_{y,n}T_{y}\to 0$
as $n\to\infty$ in $\mathbfcal{H}^{-}$, where $b^{\prime}_{y,n}=\sum_{w}b_{w,n}a_{y,w}$. For all $R>0$, there is $N>0$ such that $n>N$ implies $b_{x,n}\in(\mathbf{q}^{-1/2})^{R}\hat{\mathcal{A}}^{-}$ for all $x$. We have seen that there is $M$ depending only on $W_{\mathrm{aff}}$ such that $a_{y,w}\in(\mathbf{q}^{1/2})^{M}\hat{\mathcal{A}}^{-}$ for all $w,y$. Therefore $b^{\prime}_{y,n}\to 0$ as $n\to\infty$, because $b_{w,n}\to 0$ as $n\to\infty$. ∎
Note that Proposition 3 means in particular that $\hat{\phi}(J)\subset\mathbfcal{H}^{-}_{C^{\prime}}$.
@@
$(\phi\circ^{\dagger}(-))^{-1}(t_{1})=f_{1}(1)\sum_{w\in W_{\mathrm{aff}}}(-1)^{\ell(w)}q^{-\ell(w)}T_{w}.$
@@
If $w=1$, everything in (29) reduces to $r=d_{w}=y=1$, and we need only recall that $Q_{1,x}(\mathbf{q})=1$ for any $x\in W_{\mathrm{aff}}$. This follows from unicity of the inverse Kazhdan-Lusztig polynomials and the identity $\sum_{x\leq w}(-1)^{\ell(x)}P_{x,w}(\mathbf{q})=0$ for $w\neq 1$ [BjornerBrenti, Exercise 5.17]. ∎
@@
Let $q>1$ and let $j\neq 0$ be an element of $J$. We must show that $\tilde{\phi}(j)\neq 0$. By injectivity of $\hat{\phi}$, we have $\hat{\phi}(j)\neq 0$. By definition of the map $\hat{\phi}$, this means that there exists $q_{0}>1$, $u,s\in G^{\vee}$ such that $us=su$ with $s$ compact, and a representation $\rho$ of $\pi_{0}(Z_{G^{\vee}}(u,s))$ such that $jK(u,s,\rho,q_{0})\neq 0$. But then $jK(u,s,\rho,q)\neq 0$, $K(u,s,\rho,q)$ being a different specialization of the restriction of the same $J$-module $E(u,s,\rho)$ as for $K(u,s,\rho,q_{0})$, and $K(u,s,\rho,q)$ is also tempered. It follows that $\tilde{\phi}(j)\neq 0$. ∎
3.4 The case of $\mathrm{GL}_{n}$